Write a quadratic function f whose zeros are -11 and 3.
Answer by Edwin McCravy(20054)

Write a quadratic function f whose zeros are -11 and 3.
          x = -11;     x = 3

Get 0 on the right of each

     x + 11 = 0;   x - 3 = 0     

Multiply the left sides together and set it equal to the right sides
multiplied together:

       (x + 11)(x - 3) = 0

     x² - 3x + 11x - 33 = 0 

           x² + 8x - 33 = 0

So a quadratic function which when set equal to zero has solutions -11 and 3
is this function f:

    f(x) = x² + 8x - 33

Edwin